OpenStudy (anonymous):

The sum of two consecutive even integers is 234. What is the larger integer? 114 116 118 124

OpenStudy (anonymous):

check and see which works

OpenStudy (anonymous):

113+114=227 nope

OpenStudy (anonymous):

oh even ok 112+114 = 226 nope

OpenStudy (anonymous):

114 + 116 = ?

OpenStudy (anonymous):

230

OpenStudy (anonymous):

116 + 118 =

OpenStudy (anonymous):

234

OpenStudy (anonymous):

so now which number is larger

OpenStudy (anonymous):

118

OpenStudy (anonymous):

so that is your answer